OpenStudy (anonymous):

how to get R from P= nRT/(V-nb)-(n^2 a)/V^2 .

OpenStudy (photon336):

@sueone \[\frac{ nRT }{ (V-nb)) } - \frac{ n^2 }{ v^2 } = P\]

OpenStudy (photon336):

do you mean how do you isolate and solve for R?

OpenStudy (photon336):

\[\frac{ nRT }{ (V-nb) } = P + \frac{ n^2 }{ v^2 }\]

OpenStudy (photon336):

I guess if we want to solve for R need to first bring all of the terms over to the other side first we add the (n^2/v^2) term over and then we multiply everything by (V-nb) to get this \[(nRT) = (P + \frac{ n^2 }{ v^2 })(V-nb)\]

OpenStudy (photon336):

\[R = \frac{ (P + n^2/v^2)(V-nb) }{ nT }\]

OpenStudy (photon336):

We could say that n^2/v^2 = alpha and nb = Beta \[R = \frac{ (P + \alpha)(V-\beta) }{ nT }\] essentially this is the same thing as the ideal gas law if you notice we just added in two terms. \[PV = nRT \] \[\frac{ pV }{ nT } = R\]

OpenStudy (photon336):

if you notice it's just the same thing. by the way the real gas law, gases do interact with each-other and the walls of the container so what happens is that the real pressure is higher hence the n^2/v^2 term is added to the pressure. now the real gases themselves take up space, so the volume that the gases have in actuality to move around is less, we need to subtract from the total volume V the space the gas molecules take up which is nB.